OpenStudy (anonymous):

How many solutions are there to the following system of equations? 3x + 7y = –10 15x + 35y = –45 A. 0 B. 2 C. 1 D. infinitely many

jigglypuff314 (jigglypuff314):

3x + 7y = –10 (multiply through by 5) -> 15x + 35y = -50 15x + 35y = –45 15x + 35y = 15x + 35y right? but -50 does not equal -45 therefore no solutions

OpenStudy (anonymous):

oh my gosh i was about to put two XD im an idiot thank you so much !

jigglypuff314 (jigglypuff314):

you're welcome ;)